A 100.0 mL of 0.50 M HCl at is reacted with 125.0mL of 0.50 M NaOH in a calorimeter whose heat capacity is 500 J/C. What will the maximum temperature reached by the resulting solution if the initial temperature for the solution is 20.0°C? (Assume that the specific heat of the final solution is 4.18 J/g·°C, and that the density of the final solution is that of water.) The enthalpy change of HCl acid is neutralized by NaOH is –56 kJ/mol.
Answer #1

moles of Hal in 18000 ml of 0.50m Hl. = 0.50 imol x 198 my = 0.0 Samol inoles of Naoth in 125.0m of o.som Noott - 0.50 mo & t-0 i Enthalpy change of neutralisation reaction of lund ta & imol Noot is – 56 kJ/mol Enthalpy change heren) of neutralisatiou using equation (3) 280 = 225x 4•18 x (74 – 20) + 500 x(7-26) =) 280 = (71 -20 ) [2254498)+507 -) 2500 = (2-20) x 7490.5 =)